OpenStudy (anonymous):

Solve the following quadratic equation. x^2-9x=0

OpenStudy (anonymous):

okay so now factor the x^2-9x :)

OpenStudy (anonymous):

what would the answer be?

OpenStudy (anonymous):

you have to factor x^2-9x ....you should get (x+3)(x-3) and then you set those two equal to zero like this: x+3=0 and x-3=0 and then you solve them for x and get x=-3 and x=3

OpenStudy (lgbasallote):

careful...

OpenStudy (anonymous):

so your final answer is\[x=\pm3\]

OpenStudy (anonymous):

but 3 times 3 isn't 9x?

OpenStudy (lgbasallote):

\[x^2 - 9x = 0\] factor out \[x(x-9) = 0\] set each factor to 0 \[x = 0; \; x = 9\]

OpenStudy (anonymous):

oh i forgot there was an x :O

OpenStudy (lgbasallote):

happens to the best of us

OpenStudy (anonymous):

lol..sorry about that...thanks lgbasallote :)

OpenStudy (lgbasallote):

it was my duty as the avatar

OpenStudy (anonymous):

thanks avatar! :D